# MMEvalPro: Calibrating Multimodal Benchmarks Towards Trustworthy and Efficient Evaluation

### v1

@@ 5 Related Work
There have been several benchmarks built for evaluating LMMs, such as MMBench, MME, Seed-Bench [17, 9, 12] that assess LMMs performance from multiple fine-grained dimensions. LVLM-eHub, M3IT [27, 13] focus on the general instruction following ability. MMMU, MathVista, ScienceQA [28, 18, 19] require perception from the vision part and knowledge in the language part.
Nonetheless, critiques have been raised regarding the limitations of these existing benchmarks in effectively evaluating LMMs. PCA-Bench, MathVerse [4, 30] adopt strong LLMs such as GPT4 and GPT4-Vision to score the reasoning process of LMMs in embodied-AI and math diagram questions, in order to pinpoint cases where the LMM gets the correct answer by a fluke. Yet, using a proprietary model to conduct evaluation hinders the broader usage of the method, moreover, the evaluation result has bias itself due to the proxy model and would change over time. MMStar [6] filters out the questions that do not rely on visual information in existing multimodal benchmarks. However, it does not address the issue inherent in MCQ, where models can potentially get the correct answer without truly understanding the content. Compared with those benchmarks, MMEvalPro is more economical, easy-to-use, and calibrated for evaluating multimodal models.
@@ A.1 Data Source
MMMU [28]: The MMMU is a benchmark designed to evaluate multimodal models on massive multi-discipline tasks. The benchmark sources its questions from college examinations, quizzes, and textbooks, encompassing $30$ subjects and $183$ subfields across six core disciplines, i.e. Art & Design, Business, Science, Health & Medicine, Humanities & Social Science and Tech & Engineering. This benchmark is meticulously designed to assess models’ capabilities in handling multi-disciplinary tasks, drawing upon college-level subject knowledge.
ScienceQA [19]: The Science Question Answering (ScienceQA) is another pivotal resource, which consists of 21,208 multimodal multiple choice questions with diverse science topics. There are only $48.7\%$ questions of ScienceQA that have an image context. It is renowned for its application in multimodal tasks and features a domain diversity spanning three primary science subjects, i.e., natural, language, and social. The dataset comprises multimodal science questions that are collated from elementary and high school science curricula, ensuring a breadth of scientific inquiry and comprehension. The subjects of the questions can be categorized by Biology, Physics, Chemistry, and others.
MathVista [18]: The MathVista benchmark is developed to evaluate the reasoning ability of the multimodal models, which consists of $6,141$ examples from $31$ datasets ($28$ mathematics and IQTest, FunctionQA, PaperQA). The dataset offers exclusively mathematical and visual tasks. This source enriches our dataset with rigorous computational and analytical problems, providing a robust framework for evaluating quantitative reasoning in multimodal contexts.
@@ A.3 MMEvalPro Examples
We list more examples of different splits of MMEvalPro as shown in Figure 8, 9 and 10.
@@ A.4 Statistics of MMEvalPro
The key statistics of MMEvalPro is shown in the table 3. Figure 11and 12 display the distribution of answers and the number of options in the MMEvalPro.
Figure 11: Distribution of Answer Choices in MMEvalPro.
@@ Appendix F Comparison of Different Models
We list three cases comparing the output of GPT-4o [22], GPT-4o (non-vision) [22], InternVL-Chat [7] and LLaVA-1.5 [16]. We observed that, while both GPT-4o (non-vision) and LLaVA-1.5 could answer all the origin questions correctly, they struggled with most perception and knowledge questions. This highlights a Type-I error in current evaluation benchmarks: correctly answering a question does not necessarily indicate genuine understanding by the model. On the other hand, more advanced models like GPT-4o and InternVL-Chat demonstrate higher consistency in answering different types of questions.
Figure 18: Cases of different models’ comparison for MMEvalPro.

### v2

@@ 5 Related Work
There have been several benchmarks built for evaluating LMMs (Feng et al., 2024; Yang et al., 2024), such as MMBench, MME, Seed-Bench (Liu et al., 2023c; Fu et al., 2023; Li et al., 2023a) that assess LMMs performance from multiple fine-grained dimensions. LVLM-eHub, M3IT (Xu et al., 2023; Li et al., 2023b) focus on the general instruction following ability. MMMU, MathVista, ScienceQA (Yue et al., 2023; Lu et al., 2024, 2022) require perception from the vision part and knowledge in the language part.
Nonetheless, critiques have been raised regarding the limitations of these existing benchmarks in effectively evaluating LMMs. PCA-Bench, MathVerse (Chen et al., 2024b; Zhang et al., 2024b) adopt strong LLMs such as GPT4 and GPT4-Vision to score the reasoning process of LMMs in embodied-AI and math diagram questions, in order to pinpoint cases where the LMM gets the correct answer by a fluke. Yet, using a proprietary model to conduct evaluation hinders the broader usage of the method, moreover, the evaluation result has bias itself due to the proxy model and would change over time. MMStar (Chen et al., 2024d) filters out the questions that do not rely on visual information in existing multimodal benchmarks. However, it does not address the issue inherent in MCQ, where models can potentially get the correct answer without truly understanding the content. Compared with those benchmarks, MMEvalPro is more economical, easy-to-use, and calibrated for evaluating multimodal models.
@@ 6 Conclusion
Limitations
In order to ensure the high quality and accuracy of MMEvalPro, we employed manual annotation with human experts to construct the dataset. A certain level of human effort and expertise are necessary. To some extent, this requirement may limit the expansion efficiency of MMEvalPro.
Acknowledgments
This paper is partially supported by the National Key Research and Development Program of China with Grant No.2023YFC3341203 as well as the National Natural Science Foundation of China with Grant Numbers 61876004 and 62306014.
@@ A.1 Data Source
MMMU (Yue et al., 2023): The MMMU is a benchmark designed to evaluate multimodal models on massive multi-discipline tasks. The benchmark sources its questions from college examinations, quizzes, and textbooks, encompassing $30$ subjects and $183$ subfields across six core disciplines, i.e. Art & Design, Business, Science, Health & Medicine, Humanities & Social Science and Tech & Engineering. This benchmark is meticulously designed to assess models’ capabilities in handling multi-disciplinary tasks, drawing upon college-level subject knowledge.
ScienceQA (Lu et al., 2022): The Science Question Answering (ScienceQA) is another pivotal resource, which consists of 21,208 multimodal multiple choice questions with diverse science topics. There are only $48.7\%$ questions of ScienceQA that have an image context. It is renowned for its application in multimodal tasks and features a domain diversity spanning three primary science subjects, i.e., natural, language, and social. The dataset comprises multimodal science questions that are collated from elementary and high school science curricula, ensuring a breadth of scientific inquiry and comprehension. The subjects of the questions can be categorized by Biology, Physics, Chemistry, and others.
MathVista (Lu et al., 2024): The MathVista benchmark is developed to evaluate the reasoning ability of the multimodal models, which consists of $6,141$ examples from $31$ datasets ($28$ mathematics and IQTest, FunctionQA, PaperQA). The dataset offers exclusively mathematical and visual tasks. This source enriches our dataset with rigorous computational and analytical problems, providing a robust framework for evaluating quantitative reasoning in multimodal contexts.
@@ A.3 MMEvalPro Examples
We list examples of different splits of MMEvalPro as shown in Figure 7, 8 and 9.
@@ A.4 Statistics of MMEvalPro
The key statistics of MMEvalPro is shown in the table 4. Figure 10 display the distribution of answers and the number of options in the MMEvalPro.
@@ Appendix E Human Evaluation Guide
Figure 16: Cases of different models’ comparison for MMEvalPro.
@@ Appendix F Comparison of Different Models
We list three cases comparing the output of GPT-4o (OpenAI, 2024), GPT-4o (non-vision) (OpenAI, 2024), InternVL-Chat (Chen et al., 2023b) and LLaVA-1.5 (Liu et al., 2023b). We observed that, while both GPT-4o (non-vision) and LLaVA-1.5 could answer all the origin questions correctly, they struggled with most perception and knowledge questions. This highlights a Type-I error in current evaluation benchmarks: correctly answering a question does not necessarily indicate genuine understanding by the model. On the other hand, more advanced models like GPT-4o and InternVL-Chat demonstrate higher consistency in answering different types of questions.